# Clustered Federated Learning for Generalizable FDIA Detection in Smart Grids with Heterogeneous Data

### v1

@@ I Introduction
In smart grids, measurement devices such as Supervisory Control and Data Acquisition (SCADA) systems and Phasor Measurement Units (PMUs) are extensively deployed across geographically distributed regions to monitor real-time operational status and collect critical system data. Due to variations in physical environments, network topologies, load profiles, and regional energy structures, the collected data often exhibit pronounced spatio-temporal heterogeneity. Consequently, these data are typically non-independent and identically distributed (Non-IID) across different regions, where identical variables may follow distinct statistical distributions [1].
This distributional heterogeneity also extends to false data injection attack (FDIA) patterns, which exhibit significant geographical variability [2]. For instance, in urban areas with dense load profiles, adversaries may target voltage or power flow data to induce erroneous dispatch decisions. In contrast, in regions dominated by distributed energy resources, attackers may manipulate generator states or renewable forecasts. These spatial differences in attack vectors present substantial challenges to developing generalized detection models under conventional centralized learning paradigms.
Centralized training approaches typically require aggregating sensitive grid data, such as load curves, generation status, and market transactions from substations or regional control centers to a central server. However, such architectures pose significant privacy and security risks, including exposure to data breaches, noncompliance with regulatory frameworks, and vulnerability to single points of failure [3]. Furthermore, legal restrictions and institutional boundaries often obstruct direct data exchange among grid operators, thereby reinforcing data silos and limiting the generalization capability of learned models.
With the growing proliferation of edge computing infrastructure and sensor deployments, transmitting large-scale raw data to a central site becomes increasingly impractical due to bandwidth limitations, latency constraints, and communication overhead. These challenges underscore the urgent need for a distributed, privacy-preserving, and communication-efficient framework for training FDIA detection models in heterogeneous environments [4, 5] .
Federated Learning (FL) [6] has emerged as a promising paradigm that enables multiple clients (e.g., substations or regional controllers) to collaboratively train machine learning models without transmitting raw data. Instead, only encrypted or abstracted model updates (e.g., gradients or parameters) are exchanged, thereby preserving local data privacy. This paradigm aligns with data protection regulations such as the EU General Data Protection Regulation (GDPR) [7] and China’s Personal Information Protection Law [8], while mitigating the risks associated with data leakage [9].
@@
Although FL has demonstrated effectiveness in smart grid anomaly detection [10], several technical challenges persist:
@@
  • Non-IID Data: Heterogeneous data distributions across clients can lead to model divergence and hinder global convergence [11].
@@
  • Communication Overhead: Frequent communication between clients and the central server imposes latency and increases bandwidth usage [12].
@@ II-A Algorithmic Framework Design
At the beginning of each communication round, all clients download the latest global model parameters $\boldsymbol{w}_{0}$ from the central server. Each client then conducts multiple rounds of local training using the sampled data, resulting in an updated local model $\boldsymbol{w}_{k}$ for client $k$. Once a sufficient number of client updates $\{\boldsymbol{w}_{k}\}$ are collected, the server aggregates these models to produce a new global model ${\boldsymbol{w}_{0}}^{{}^{\prime}}$, which is then broadcast to all clients for the next iteration.
@@
      Calculate the distance between all other samples and the existing $J$ cluster centers:$D_{i1},D_{i2},\ldots,D_{ij}\\ \left(i=1,2,\ldots,n_{k}\right)$.
@@
To streamline the algorithmic design, we impose two simplifying assumptions on the local client dataset. First, the sample size of each client is required to exceed 100, ensuring sufficient statistical support for clustering. Second, the distance between cluster centers in the optimal classification scheme must exceed $50\%$ of the average pairwise distance between samples prior to clustering. This condition guarantees adequate separability among clusters.
@@
  • Small clients obtained by clustering($j$ category, the sample size of each category is $n_{kj}$)
@@ III-A Experimental Setup
FDIA detection is performed using the Rec-AD model we proposed in reference , integrated within the FedClusAvg framework. The GPU configuration is the same as Rec-AD in our experiment.
@@ III-C Generalization Analysis
Client heterogeneity FedAvg+ Accuracy FedClusAvg+ Accuracy FedAvg+ Accuracy FedClusAvg+ Accuracy
IEEE300-Bus IEEE300-Bus Provincial power grid Provincial power grid
@@ III-D Communication efficiency analysis
As illustrated in Fig. 6, empirical results on the IEEE 118-bus dataset demonstrate that under WiFi conditions (10 Mbps), FedClusAvg incurs an average communication delay of approximately 320 seconds per round, while FedClusAvg+ reduces this delay to around 270 seconds—yielding a reduction of roughly 15.6%. Under 5G conditions (100 Mbps), the advantage becomes more pronounced: FedClusAvg+ achieves a per-round delay of 85 seconds, compared to 100 seconds for FedClusAvg. These findings highlight the superior bandwidth utilization and network adaptability of the hierarchical FedClusAvg+ design.
Moreover, FedClusAvg+ supports enhanced concurrency by enabling multiple sub-servers to concurrently aggregate updates from different subsets of clients. This mitigates the congestion typically observed when numerous clients simultaneously communicate with a single central server. Consequently, FedClusAvg+ not only improves communication efficiency and reduces latency but also provides a more scalable foundation for deployment in large-scale federated environments.
@@ IV Conclusion
We proposed an enhanced federated learning framework based on the FedClusAvg algorithm to address key challenges in False Data Injection Attack (FDIA) detection for smart grids. The framework tackles the Non-IID nature of distributed measurements and preserves data privacy through a hierarchical architecture that integrates clients, sub-servers, and a central server.
By incorporating clustered sampling, deviation-based weighted aggregation, and multi-tier communication, the proposed method improves model generalization, detection accuracy, and training efficiency. It significantly reduces communication rounds and bandwidth consumption while enhancing robustness against regional misclassification and adversarial exploitation of data heterogeneity. Experimental evaluations on the IEEE 118-bus and 300-bus systems demonstrate that the proposed algorithm consistently outperforms classical methods such as FedAvg, FedProx, and FedNova in terms of accuracy, recall, and communication efficiency. Moreover, the framework ensures privacy compliance and reduces the risk of data breaches, making it suitable for real-world deployment in large-scale, heterogeneous, and security-critical power systems. In summary, the proposed approach advances the development of secure, scalable, and privacy-preserving FDIA detection solutions in modern smart grids.

### v2

@@ I Introduction
As an important cyber-physical system (CPS), smart grid is highly vulnerable to cyber attacks [1]. In modern smart grid infrastructures, measurement devices such as Supervisory Control and Data Acquisition (SCADA) systems and Phasor Measurement Units (PMUs) are extensively deployed across geographically distributed regions to monitor real-time operational status and collect critical system data. Due to variations in physical environments, network topologies, load profiles, and regional energy mixes, the acquired data is often highly heterogeneous, exhibiting pronounced spatio-temporal variability and non-independent and identically distributed (Non-IID) characteristics across regions, where even identically-named variables may follow distinct distributions [2, 3, 4].
This statistical heterogeneity further extends to cyberattack behaviors, particularly in the context of false data injection attacks (FDIAs), where attack vectors demonstrate spatial patterns. For instance, adversaries in urban areas may compromise power flow data to induce incorrect economic dispatch decisions, while attackers in regions rich in renewable energy may manipulate forecast signals or distributed generator states [5, 6]. These non-uniform attack patterns pose significant challenges to traditional centralized machine learning-based detection frameworks, which assume homogenous training data.
Centralized FDIA detection models typically require aggregating sensitive grid data—such as substation measurements, load consumption profiles, and generation forecasts—into a centralized data center. However, such architectures face growing scrutiny due to the risks of privacy leakage [7], compliance violations (e.g., GDPR [8]), and exposure to cyber-physical attacks or system-wide failures [9, 10]. Moreover, institutional and jurisdictional boundaries often inhibit direct data sharing among regional grid operators, reinforcing data silos and weakening the global generalizability of trained models. The increasing deployment of edge computing platforms and the rising volume of sensor data also make centralized processing cost-prohibitive, suffering from latency bottlenecks and bandwidth limitations [11]. These practical concerns underscore the urgent demand for privacy-preserving, bandwidth-efficient, and generalizable frameworks for FDIA detection under real-world heterogeneous conditions.
Federated Learning (FL) has emerged as a transformative paradigm that enables decentralized model training across distributed clients without exchanging raw data [12]. Instead, clients periodically transmit encrypted or abstracted model updates-such as gradients or model weights-to a server for aggregation [13, 14]. This architecture not only preserves local data privacy and aligns with regional regulations, but also supports scalability in geographically distributed energy systems [15].
To effectively detect FDIA in smart grids, researchers have developed several privacy-preserving FL approaches. These include implementations using the Paillier cryptosystem and Local Differential Privacy (LDP) techniques [16]. Additionally, decentralized FL frameworks have been specifically designed to identify false data injection attacks targeting solar PV dc/dc and dc/ac converters [17], as well as to detect fraudulent behavior among advanced metering infrastructure consumers who compromise their meters [18]. To enhance the robustness of FL-based machine learning models against adversarial attacks during both training and inference phases, researchers have explored novel FL-based false data detection methods incorporating Explainable Artificial Intelligence [19]. Furthermore, to address challenges posed by unknown system parameters and limited decentralized datasets with strategic data owners, incentive-based edge FL mechanisms have been developed [20]. Several advanced frameworks have emerged to address specific smart grid security challenges. These include communication-efficient and privacy-preserving FL schemes for secure energy data sharing through edge-cloud collaboration [12], and collaborative learning frameworks utilizing vertical FL for detecting false data injection attacks [21]. To protect against inference attacks from potentially dishonest aggregators who might extract information about clients’ training data from model parameters, efficient cross-silo FL schemes with enhanced privacy preservation have been developed [22]. Moreover, researchers have developed FL schemes that exploits local correlations between connected power buses through graph neural networks while capturing temporal patterns using long short-term memory layers [23]. Nevertheless, FL suffers from inherent performance degradation under highly non-IID settings [24]. To address this, Clustered Federated Learning (CFL) strategies have been proposed, which group clients based on statistical similarity or attack vector profiles, and train models within each cluster separately [25, 26, 27]. This clustering not only improves local model relevance but also enhances robustness to adversarial manipulations.
@@
Although FL has demonstrated effectiveness in smart grid anomaly detection, several technical challenges persist [28]:
@@
  • Non-IID Data: Heterogeneous data distributions across clients can lead to model divergence and hinder global convergence.
@@
  • Communication Overhead: Frequent communication between clients and the central server imposes latency and increases bandwidth usage.
@@ II-A Algorithmic Framework Design
At the beginning of each communication round, all clients download the latest global model parameters $\boldsymbol{w}_{0}$ from the central server. Each client then conducts multiple rounds of local training using the sampled data, resulting in an updated local model $\boldsymbol{w}_{k}$ for client $k$. Once a sufficient number of client updates $\boldsymbol{w}_{k}$ are collected, the server aggregates these models to produce a new global model ${\boldsymbol{w}_{0}}^{{}^{\prime}}$, which is then broadcast to all clients for the next iteration.
@@
[…]
@@
[…]
@@
[…]
@@ III-A Experimental Setup
FDIA detection is performed using the Rec-AD model we proposed in reference [31], integrated within the FedClusAvg framework. The GPU configuration is the same as Rec-AD in our experiment.
@@ III-C Generalization Analysis
Heterogeneity FedAvg+ FedClusAvg+ FedAvg+ FedClusAvg+
300-Bus 300-Bus Provincial Provincial
@@ III-D Communication efficiency analysis
As illustrated in Fig. 6, empirical results on the IEEE 118-bus dataset demonstrate that under WiFi conditions (10 Mbps), FedClusAvg incurs an average communication delay of approximately 320 seconds per round, while FedClusAvg+ reduces this delay to around 270 seconds—yielding a reduction of roughly 15.6%. Under 5G conditions (100 Mbps), the advantage becomes more pronounced: FedClusAvg+ achieves a per-round delay of 85 seconds, compared to 100 seconds for FedClusAvg. These findings highlight the superior bandwidth utilization and network adaptability of the hierarchical FedClusAvg+ design. Moreover, FedClusAvg+ supports enhanced concurrency by enabling multiple sub-servers to concurrently aggregate updates from different subsets of clients. This mitigates the congestion typically observed when numerous clients simultaneously communicate with a single central server. Consequently, FedClusAvg+ not only improves communication efficiency and reduces latency but also provides a more scalable foundation for deployment in large-scale federated environments.
@@ IV Conclusion
We proposed an enhanced federated learning framework based on the FedClusAvg algorithm to address key challenges in False Data Injection Attack (FDIA) detection for smart grids. The framework tackles the Non-IID nature of distributed measurements and preserves data privacy through a hierarchical architecture that integrates clients, sub-servers, and a central server. By incorporating clustered sampling, deviation-based weighted aggregation, and multi-tier communication, the proposed method improves model generalization, detection accuracy, and training efficiency. It significantly reduces communication rounds and bandwidth consumption while enhancing robustness against regional misclassification and adversarial exploitation of data heterogeneity. Experimental evaluations on the IEEE 118-bus and 300-bus systems demonstrate that the proposed algorithm consistently outperforms classical methods such as FedAvg, FedProx, and FedNova in terms of accuracy, recall, and communication efficiency. Moreover, the framework ensures privacy compliance and reduces the risk of data breaches, making it suitable for real-world deployment in large-scale, heterogeneous, and security-critical power systems. In summary, the proposed approach advances the development of secure, scalable, and privacy-preserving FDIA detection solutions in modern smart grids.